Hello! Can someone help me with the problem below. not sure if my answers are correct:
Assume that you have surveyed 1068 randomly selected adult Americans and 279 have indicated that they would vote for Andrew Mason if he were to run for Congress. Construct a 95% confidence interval for the percentage of voters who would vote for Mason. State the margin of error.
After computing, my margin of error is = .0263 and my CI is = .2349 < P < .2875
is this correct? Thank you!

Answer by stanbon(57387)

Assume that you have surveyed 1068 randomly selected adult Americans and 279 have indicated that they would vote for Andrew Mason if he were to run for Congress. Construct a 95% confidence interval for the percentage of voters who would vote for Mason. State the margin of error.
-----
phat = 279/1068 = 0.2612
ME = 1.96*sqrt[0.26*0.74/1068] = 0.0263
95% CI : 0.2612-0.0263 < p < 0.2612+0.0263

After computing, my margin of error is = .0263 and my CI is = .2349 < P < .2875
is this correct?.
-----------------------
You are correct.
-------------------
Cheers,
Stan H.